OpenStudy (anonymous):

This month a company's cost to manufacture 50 floopers was 8000 dollars. Their marginal cost for the same number of floopers was 700 dollars. Use the above information to estimate their cost to produce 53 floopers. Cost is approximately ??

OpenStudy (anonymous):

Total Cost = Fixed Cost + Marginal Cost Where Fixed Cost = Cost not depending on unit produced (Like cost of setting up a factory) Marginal Cost = Cost proportional to number of unit produced Since for 50 units MC = $700 => for 1 unit MC = 700 / 50 = $14 For 53 Unit MC = 53 * 14 = $742 Again, Fixed Cost = 8000 - 700 = $7300 Hence Total Cost for 53 unit = 7300 + 742= $8042